Session: Feb 2000 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Cattani at Wed Feb 16 18:43:21 MET 2000) Success

Effelsberg (A. Kraus at Thu Feb 17 08:31:05 MET 2000) Minor failures
Due to snowfall the scans between UT1457-1533 and UT1603-1736 had been lost.

Torun (K.Borkowski at Thu Feb 17 09:50:31 MET 2000) Success
No known problems.

Onsala (A. Polatidis at Thu Feb 17 12:33:21 MET 2000) Success

Noto (C.Trigilio at Thu Feb 17 13:03:54 MET 2000) Minor failures
Data OK from start to 7:54. VLBA FORMATTER FAILURE !!! Switch to MK4 Formatter. Schedule and procedures re-created considering the new configuration (VLBA4 rack, VLBA recorder in drudg). Re-start of observation at 12:04 UT. NO DATA RECORDED FROM 7:54 TO 12:04 UT.

Westerbork (Tony Foley at Thu Feb 17 17:48:17 MET 2000) Minor failures
No known problems. Full tied array used. During this experiment workmen were replacing part of the roof of Westerbork. Full tied array used.

Effelsberg (dgraham at Mon Feb 21 08:33:13 MET 2000) Minor failures
due to computer disc failure, log file was lost

Jodrell (Alastair Gunn at Wed Feb 23 14:35:54 MET 2000) Minor failures
Elevation drive fail at 11:06 UT, back on at 11:20 UT. Telescope was sent early to some sources due to insufficient slew times provided by PI. No other known problems.

Calibration (Alastair Gunn at Wed Apr 19 13:01:58 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/ev009.antab for details